OpenStudy (oblivion20):

Which pair of ratios is proportional? A. 6/21and 2/42 B. 25/100 and 50/150 C. 18/24 and 6/12 D. 7/14 and 20/40 I will Fan and give medal!

OpenStudy (igreen):

Divide them all see which one's are equal to each other.

OpenStudy (igreen):

6 / 21 = ? 2 / 42 = ?

OpenStudy (oblivion20):

6/21 = 3.5? and 2/42 =21?

OpenStudy (oblivion20):

so not proportional?

OpenStudy (igreen):

Actually you divided 21 / 6 and 42 / 2..but it doesn't matter, yes, they're not proportional, now try option B.

OpenStudy (oblivion20):

ok

OpenStudy (oblivion20):

25/100= 0.25 and 50/150= 0.33

OpenStudy (oblivion20):

so still not proportional?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

keep going

OpenStudy (oblivion20):

ok thank both of you i think i got it now :)

OpenStudy (oblivion20):

but just to make sure

OpenStudy (oblivion20):

18/24=0.75 and 6/12=0.5?

OpenStudy (anonymous):

yep, the calculator is safest route, but i would try to reduce the easiest one in my head, for example, 6/12 you can look at and right away know its 1/2 (and 1 is half the size of 2, numerator is half size of denominator), so i look at the other one 18/24 and say whats a number half of 24 since i want the numerator to be half the size of the denominator, ok thats 12, since the numerator is 18 which is larger than 12 then they are obviously not proportional and 18/24 is larger than 6/12

OpenStudy (igreen):

Yes..now try D..

OpenStudy (igreen):

Yeah, there are multiple ways to do this..you can even cross multiply.

OpenStudy (oblivion20):

7/14=0.5 and 20/40=0.5

OpenStudy (oblivion20):

so they are proportional?

OpenStudy (igreen):

Yep! So D is your answer.

OpenStudy (oblivion20):

ok thank you